OpenStudy (anonymous):

1.0 kg of water is warmed from room temperature (25.0°C) to boiling, then half of its initial volume is vaporized. How many joules of thermal energy must the stove provide to do this? Cwater = 2.26×106 J/(kg·°C); Hv, water = 2.26×106 J/kg

OpenStudy (domebotnos):

Welcome to openstudy!

OpenStudy (shamim):

Heat needed to become 100 degree celcius of water frm 25 degree celcius Q=ms*del theta Q=1*4200*(100-25)=?

OpenStudy (shamim):

Right?

OpenStudy (anonymous):

don't we need to apply the Q=mHv equation?

OpenStudy (anonymous):

or is it the same thing?

OpenStudy (anonymous):

where did you get the 4200 from?

OpenStudy (shamim):

Now heat needed to become vapour Q=mLv Q=0.5*2.26*10^6=?

OpenStudy (shamim):

U hv to add those 2 Q

OpenStudy (shamim):

U will get total heat needed

OpenStudy (anonymous):

thank u!

OpenStudy (shamim):

Anyway s is specific heat of watet

OpenStudy (shamim):

Welcome!!!"

OpenStudy (shamim):

I think u got it!!!!

OpenStudy (shamim):

Medal??????!!!!

OpenStudy (anonymous):

after i calculate the Q, wont i have to cut my result in half because half is vaporized?

OpenStudy (shamim):

No!!!

OpenStudy (shamim):

I used m=0.5kg during vaporization

OpenStudy (shamim):

U know total mass 1 kg water become 100 degree celcius of water

OpenStudy (anonymous):

yes

OpenStudy (shamim):

But 0.5 kg water become vapour frm 100 degree celcius

OpenStudy (anonymous):

right! yay, i got it! thank you so much! u were extremely helpful!

OpenStudy (shamim):

Glad to help u!!!!